Container A is half full of water. Container B will hold 9 cups of water. When the contents of container A is poured into container B, container B is 2/3 full. How much will container A hold?
Answer by nerdybill(6948)

Container A is half full of water. Container B will hold 9 cups of water. When the contents of container A is poured into container B, container B is 2/3 full. How much will container A hold?
.
Let x = amount Container A can hold
then
(1/2)x = (2/3)(9)
(1/2)x = (2)(3)
(1/2)x = 6
x = 12
.
Container A can hold 12 cups